# Nova: Generative Language Models for Assembly Code with Hierarchical Attention and Contrastive Learning

### v4

@@ 3.3 Binary Code Similarity Detection
Training Data To compare Nova with existing works on BCSD fairly Wang et al. (2022); Su et al. (2024), we use BinaryCorp-3M Wang et al. (2022) as the fine-tuning data for BCSD, which contains the O0 and O3 assembly of 224,606 functions.
@@
Test Data Following existing work Su et al. (2024); Xu et al. (2023a), we use real-world benchmarks, Binutils, Curl, ImageMagick, SQLite, OpenSSL, and Putty, as the test benchmarks, which are nonexistent in the training data.
@@
Baselines Nova is compared with jTrans Wang et al. (2022), DiEmph Xu et al. (2023a) and CodeArt Su et al. (2024). jTrans is a Transformer Vaswani et al. (2023) encoder trained on binaries with masked token prediction and jump target prediction tasks. DiEmph uses an instruction deemphasis technique to prevent the model from learning instruction distribution biases introduced by compilers. CodeArt proposes a regularized attention mask for encoder models to capture instructional semantics and data dependencies.
@@
Evaluation We randomly sample $K$ source code functions from each project, compile them into binaries with O0 and O3 optimization flags, and disassemble them into X86-64 assemblies. BCSD techniques encode these assemblies into embeddings (Nova uses the average embeddings of all the “[INST]” tokens in an assembly as its embedding). Then each O0 assembly is used as the query to calculate their similarity with the $K$ O3 candidate assemblies (using cosine similarity). Metric Recall@1 is the ratio of queries for which the candidate from the same source code has the highest similarity among all the candidates.
Appendix A.3 contains additional details such as training hyper-parameters.
@@ 4.1.1 Comparison with SOTA Techniques
Table 2 shows the Pass@1 of the decompiled code from assemblies on the HumanEval-Decompile benchmark. The results are grouped by optimization level (i.e., the benchmark contains 164 assemblies of each optimization level to decompile), and the average is also reported.
Optimization O0 O1 O2 O3 Avg.
GPT-3.5 6.80 5.64 4.36 3.93 5.18
GPT-4 17.77 15.12 12.65 11.25 14.20
LLM4Decompile-1B 12.26 7.22 8.38 7.96 8.96
Nova-1B 31.19 17.29 18.72 15.58 22.09
LLM4Decompile-6B 23.01 15.95 16.95 14.79 17.68
Nova-6B 42.07 28.04 25.00 22.56 29.42
@@
Table 2: Pass@1 on HumanEval-Decompile.
Optimization O0 O1 O2 O3 Avg.
GPT-3.5 8.95 7.77 5.93 5.12 6.94
GPT-4 25.64 20.65 18.70 18.03 20.76
LLM4Decompile-1B 17.95 12.05 13.90 12.51 14.10
Nova-1B 41.11 29.81 31.18 26.24 32.09
LLM4Decompile-6B 33.77 24.25 23.94 23.81 26.44
Nova-6B 51.06 37.83 35.79 34.63 39.83
Table 3: Pass@10 on HumanEval-Decompile.
@@
Overall, Nova’s Pass@1 is higher than all SOTA binary decompilation techniques and general LLMs GPT-4 and GPT-3.5, which are orders of magnitude larger than Nova. Specifically, for each optimization level, Nova consistently decompiles more assemblies into source code correctly than LLM4Decompile, GPT-3.5, and GPT-4. With the same model size, Nova-1B outperforms LLM4Decompile-1B by 146.54%, i.e., averaged Pass@1 of 22.09% versus 8.96%. Nova-6B outperforms LLM4Decompile-6B by 66.40%: the averaged Pass@1 is 29.42% versus 17.68%.
Table 3 shows that Nova still outperforms SOTA techniques with a significant margin under the measurement of Pass@10. Examples of Nova’s correct decompilation are provided in Appendix A.4.
4.1.2 Ablation Study
@@
We conduct an ablation study by comparing Nova-1B with the following models:
@@
  • Nova-CL-HA: Removing contrastive learning and hierarchical self-attention. This is simply training DeepSeek-Coder-1.3B on the assembly corpus using language modeling.
@@
  • Nova-HA: Removing the hierarchical self-attention, training DeepSeek-Coder-1.3B on the assembly corpus using both the language modeling and contrastive learning objectives.
Nova-CL-HA can be viewed as our reproduction (retrain) of LLM4Decompile-1B.
Optimization O0 O1 O2 O3 Avg.
LLM4Decompile-1B 17.95 12.05 13.90 12.51 14.10
Nova-CL-HA 17.80 13.32 13.26 10.03 13.60
Nova-HA 25.12 15.64 16.07 12.71 17.39
Nova 31.19 17.29 18.72 15.58 22.09
Table 4: Ablation study of Nova-1B (Pass@1).
@@
Table 4 shows the results of the ablation study, reported by the Pass@1 on HumanEval-Decompile. Nova-CL-HA shows comparable Pass@1, which we considered as variance in reproducing the same approach. With additional contrastive learning objectives, Nova-HA improves the Pass@1 on all optimization levels over Nova-CL-HA, showing a 27.87% higher averaged Pass@1. Further applying the hierarchical self-attention boosts the overall Pass@1 from 17.39% to 22.09%.
@@
Table 5: Recall@1 on benchmarks with $K=50$.
@@
Table 6: Recall@1 on benchmarks with $K=100$.
@@
Table 7: Recall@1 on benchmarks with $K=200$.
@@ 4.2 Binary Code Similarity Detection
4.2.1 Comparison with SOTA Techniques
Tables 8,  8,  8 and  8 show the Recall@1 of Nova and existing BCSD techniques with pool size $K$ of 50, 100, 200 and 500 on the six benchmarks. Underlined numbers indicates the best in each benchmark, while wavey underlined numbers denote the tied best (we only mark Nova-1B for clearer illustration).
Overall, Tables 888 and 8 show that on average, Nova-1B and Nova-6B achieve the highest Recall@1 (in bold) under all four settings of $K$. Nova-6B further outperforms Nova-1B and achieves the highest averaged Recall@1 under all four settings, ranking the ground-truth of 5%, 2%, 4%, and 3% more queries the most similar correspondingly compared to CodeArt.
Nova-1B consistently outperforms existing techniques with higher Recall@1 when $K$ is 50, 100, and 200, meaning it correctly ranks ground-truth of 3%, 1%, and 2% more queries as the most similar. Under the setting of $K=500$, Nova-1B ties with CodeArt with the same highest Recall@1. When looking into each individual benchmark, Nova-1B always wins on the most benchmarks under different settings of pool size $K$. For instance, Nova-1B wins on four benchmarks while DiEmph only wins on SQLite when $K=50$.
$K$ Nova-CL-HA Nova-HA Nova
50 0.81 0.83 (+0.02) 0.84 (+0.01)
100 0.76 0.78 (+0.02) 0.78
200 0.70 0.70 0.73 (+0.03)
500 0.60 0.62 (+0.02) 0.64 (+0.02)
Table 9: Ablation study of Nova-1B (Recall@1)
4.2.2 Ablation Study
Table 9 shows the averaged Recall@1 of Nova-CL-HA, Nova-HA (same as in Section 4.1.2), and Nova-1B under four pool size settings. With contrastive learning objectives, Nova-HA improves Nova-CL-HA under three settings ($K=50,100,200$) with 2% higher Recall@1. With hierarchical attention, Nova further outperforms Nova-HA under three settings ($K=50,200,500$). Detailed ablation study results on each benchmark are provided in Appendix A.5.
4.3 Analytic Experiments
4.3.1 How are Nova’s embeddings better?
@@
Figure 5 (b) shows that Nova encodes assemblies into clusters of functionalities. The assemblies for the same functionality (i.e., the same task) are encoded closer to each other than Nova-CL-HA does in Figure 5 (b). The results show that our hierarchical attention and contrastive learning techniques effectively group codes of similar functionalities together for better assembly foundation models. Embedding of Nova-HA and additional quantitative analysis are shown in Appendix A.6.
Figure 5: PCA of embeddings calculated by Nova-CL-HA and Nova, for HumanEval-Decompile assemblies.
4.3.2 What does hierarchical attention learn?
We conduct quantitative analysis on the attention weights produced by different models.
Figure 6: Quantitative analysis of attention weights.
Entropy Figure 6 (a) shows the entropy of attention-weight distributions in each layer. We separate the attention heads as standard attention (green line) and hierarchical attention (red line), since Nova applies hierarchical attention to half of the attention heads in each layer (Section 3.1).
Nova’s hierarchical attention heads produce significantly lower entropy, suggesting its attention layer is more confident in learning specific relationships than the other models’ attention layers. The standard attention heads in Nova show patterns similar to those of Nova-CL-HA and Nova-HA, allowing Nova learning the standard attention to capture the “soft” relationship between possible tokens pairs. The result suggests that hierarchical attention complements standard attention with additional knowledge.
[INST] Token Figure 6 (b) shows the attention weights paid to the “[INST]” tokens. Nova’s hierarchical attention heads pay more attention to the “[INST]” tokens than standard attention, which may be because these “[INST]” tokens contain instruction summary and long dependencies and thus are more informative. Additional analysis and examples are given in Appendix A.7.
@@ 5.1 Binary Models
Machine learning models are widely used in binary program analysis tasks. However, these models are typically designed for specific downstream tasks such as binary code similarity detection Pei et al. (2020); Xu et al. (2023a); Wang et al. (2022), variable name prediction Chen et al. (2022c); Xu et al. (2023b); Zhang et al. (2021), binary code type inference Pei et al. (2021), and are built from scratch.
Recent techniques have started to pre-train foundation LLMs for binaries. CodeArt Su et al. (2024) pre-trains encoder-style LLMs with a regularized attention design to better encode assembly code semantics, showing good accuracy on binary code understanding tasks (e.g., binary code similarity detection and malware family classification). SLaDe Armengol-Estapé et al. (2024) trains BART Lewis et al. (2019) models on assembly, and LLM4Decompile Tan et al. (2024) trains DeepSeek-Coder with assembly for binary code decompilation. However, CodeArt does not generalize to generation tasks due to its encoder architecture. SLaDe and LLM4Decompile are limited in performance due to a lack of special designs for assembly. In contrast, Nova addresses both limitations, by using the proposed hierarchical attention and contrastive learning objectives, outperforming existing techniques on both understanding (binary code similarity detection) and generation (binary code decompilation) tasks.
@@ 5.2 Large Source-Code Models
LLMs demonstrate promising results on many code-related tasks, such as code generation Chen et al. (2022a); Liu et al. (2023); Chen et al. (2023); Le et al. (2022); Yue et al. (2021); Chen et al. (2021); Nijkamp et al. (2022); Fried et al. (2023); Rozière et al. (2023); Guo et al. (2024), bug fixing Jiang et al. (2023); Xia et al. (2023) and vulnerability fixing Wu et al. (2023); Steenhoek et al. (2023); He and Vechev (2023). The advances in using LLMs are attributed to the knowledge learned from massive source code and natural language text in their training datasets Touvron et al. (2023); OpenAI (2023). Nova is designed and trained for assembly, which has unique challenges such as low information density and diverse optimization.

### v5

@@ 3.3 Fine-Tuning for Binary Code Similarity Detection
Training Data: To compare Nova with existing works on BCSD fairly (Wang et al., 2022; Su et al., 2024), we use the same dataset, BinaryCorp-3M (Wang et al., 2022), as the fine-tuning data for BCSD, which contains the O0 and O3 assembly of 224,606 functions.
@@
Test Data: Following existing work (Su et al., 2024; Xu et al., 2023a), we use real-world benchmarks, Binutils, Curl, ImageMagick, SQLite, OpenSSL, and Putty, as the test benchmarks, which are nonexistent in the training data.
@@
Baselines: Nova is compared with jTrans (Wang et al., 2022), DiEmph (Xu et al., 2023a) and CodeArt (Su et al., 2024). jTrans is a Transformer (Vaswani et al., 2023) encoder trained on binaries with masked token prediction and jump target prediction tasks. DiEmph uses an instruction deemphasis technique to prevent the model from learning instruction distribution biases introduced by compilers. CodeArt proposes a regularized attention mask for encoder models to capture instructional semantics and data dependencies.
@@
Evaluation: We randomly sample $K$ source code functions from each project, compile them into binaries with O0 and O3 optimization flags, and disassemble them into X86-64 assemblies. BCSD techniques encode these assemblies into embeddings (Nova uses the average last-layer hidden states of all the “[INST]” tokens in an assembly as its embedding). Then each O0 assembly is used as the query to calculate their similarity with the $K$ O3 candidate assemblies (using cosine similarity). Metric Recall@1 is the ratio of queries for which the candidate from the same source code has the highest similarity among all the candidates.
@@ 4.1 Binary Code Decompilation
Comparison with SOTA Techniques: Table 2 shows the Pass@1 and Pass@10 of the decompiled code from assemblies on HumanEval-Decompile. The results are grouped by optimization level (i.e., the benchmark contains 164 assemblies of each optimization level), and the average is also reported.
@@
Table 2: Nova’s Pass@K and comparison with existing techniques on HumanEval-Decompile.
Techniques Pass@1 Pass@10
O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
CodeLlama-7B 6.95 3.81 4.54 3.78 4.77 8.53 5.97 7.34 5.17 6.75
StarCoder2-7B 6.31 4.33 5.64 5.95 5.56 8.77 5.18 6.09 7.17 6.80
DeepSeekCoder-7B 9.63 7.56 7.41 6.68 7.82 13.60 11.38 11.52 9.44 11.49
Qwen-2.5-Coder-7B 4.76 5.79 5.58 5.27 5.35 6.34 7.69 6.56 5.79 6.60
LLMCompiler-7B 5.95 5.85 5.55 5.82 5.79 7.01 7.31 7.47 7.01 7.20
GPT-3.5-Turbo 7.41 6.13 4.33 3.90 5.44 9.56 8.38 6.23 5.12 7.32
GPT-4o 21.34 18.29 14.48 13.05 16.79 29.94 26.74 21.42 19.88 24.50
LLM4Decompile-1B 15.30 8.26 9.36 8.38 10.33 21.79 15.23 16.17 13.70 16.72
Nova-1B 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
LLM4Decompile-6B 29.97 19.05 20.46 18.32 21.95 40.40 27.75 28.85 28.51 31.38
Nova-6B 48.78 30.58 30.85 27.23 34.36 57.47 47.45 43.03 39.68 46.91
@@
Overall, Nova’s Pass@1 and Pass@10 are higher than all SOTA binary decompilation techniques and general LLMs with even smaller model sizes. Specifically, for each optimization level, Nova consistently decompiles more assemblies into source code correctly than the rest of the compared techniques. Note that Meta LLMCompiler is mainly designed for LLVM IR optimization, and thus is still incapable of assembly code decompilation.
With the same model size, Nova-1B outperforms LLM4Decompile-1B with a 14.84% higher averaged Pass@1, and a 21.58% higher Pass@10. Nova-6B outperforms LLM4Decompile-6B with a 12.41% higher averaged Pass@1, and a 15.53% higher Pass@10. When compared with GPT-4o, an order of magnitude larger model, Nova-1B produces an 8.38% higher Pass@1 and 13.80% higher Pass@10. Examples of Nova’s correct decompilation are provided in Appendix A.4.
Comparison with Techniques Handling Long Input: Nova’s hierarchical attention design targets to address the low information density and long input challenge of assembly code. There are other techniques that handles long input challenges in text and source code, with Granite-3B-Code-Base-128K and LongCoder being the most related ones. Granite trains LLM on repository-level long inputs, which is an orthogonal approach with Nova’s approach (hierarchical attention and contrastive learning). We train Granite-3B-Code-128K with Nova’s approach, and Table 3 shows that Nova’s approach brings improvement to Granite over standard fine-tuning even if it has already been trained with long code data.
Table 3: Nova’s approach brings improvement to LLM that has been trained with long input data.
Techniques Pass@1 Pass@10
O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
Granite (3B-Code-128K) 5.91 3.78 5.09 5.52 5.08 8.19 5.16 6.56 7.15 6.76
Granite + Standard Fine-Tuning 20.88 13.54 11.37 10.09 13.97 30.05 19.77 18.31 15.77 20.98
Granite + Nova’s Fine-Tuning 31.04 14.57 14.70 13.66 18.49 39.57 21.23 21.77 19.82 25.60
Table 4: Nova’s hierarchical attention is more effective on assembly code.
Techniques Pass@1 Pass@10
O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
DeepSeekCoder + LongCoder’s Attention 34.59 19.07 19.72 17.34 22.68 42.19 32.37 32.86 29.04 34.12
DeepSeekCoder + Nova’s Attention 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
LongCoder combines window attention and global attention to learn long code input. We compare LongCoder’s attention design with Nova’s hierarchical attention design using DeepSeekCoder-1.3B as the backbone. Table 4 shows that Nova’s hierarchical attention is more effective in learning assembly code. Nova’s attention design considers the instruction-level local semantics and dependencies between different instructions, which fits better than fix-sized window attention to assembly code.
@@
Ablation Study: We conduct an ablation study by comparing Nova-1B with the following models to show the effectiveness of contrastive learning objectives and hierarchical attention design:
@@
  • Nova-CL-HA: Removing contrastive learning and hierarchical self-attention. This is simply training DeepSeekCoder-1.3B on the assembly corpus using language modeling. This can be viewed as our reproduction (retrain) of LLM4Decompile-1B using the same amount of data.
@@
  • Nova-HA: Removing the hierarchical self-attention, training DeepSeekCoder-1.3B on the assembly corpus using both the language modeling and contrastive learning objectives.
Table 5: Ablation study of Nova-1B on HumanEval-Decompile.
Techniques Pass@1 Pass@10
O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
LLM4Decompile-1B 15.30 8.26 9.36 8.38 10.33 21.79 15.23 16.17 13.70 16.72
Nova-CL-HA 20.73 16.16 15.03 11.19 15.78 33.55 28.12 26.96 21.01 27.41
Nova-HA 30.58 19.88 20.58 16.40 21.86 44.75 33.13 33.31 29.82 35.25
Nova 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
@@
Table 5 shows the results of the ablation study. Nova-CL-HA produces an average Pass@1 of 15.78% and Pass@10 of 27.41%. With additional contrastive learning objectives, Nova-HA improves the Pass@1 on all optimization levels over Nova-CL-HA, showing a higher averaged Pass@1 and Pass@10. Further applying the hierarchical self-attention on Nova-HA boosts the overall Pass@1 from 21.86% to 25.17%, and Pass@10 from 35.25% to 38.30%.
Table 6: Recall@1 on BCSD with $K=50$
@@
Table 7: Recall@1 on BCSD with $K=100$
@@
Table 8: Recall@1 on BCSD with $K=200$
@@
Table 9: Recall@1 on BCSD with $K=500$
@@ 4.2 Binary Code Similarity Detection
Tables 9,  9,  9 and  9 show the Recall@1 of Nova and existing BCSD techniques with pool size $K$ of 50, 100, 200 and 500 on the six benchmarks. Underline indicates the best in each benchmark, while wave denotes the tied best (we only mark Nova-1B for clearer illustration).
Overall, Tables 9,  9,  9 and 9 show that on average, Nova-1B and Nova-6B achieve the highest Recall@1 (in bold) under all four settings of $K$. Nova-6B further outperforms Nova-1B and achieves the highest averaged Recall@1 under all four settings, ranking the ground-truth of 5%, 2%, 4%, and 3% more queries the most similar correspondingly compared to CodeArt. Nova-1B consistently outperforms existing techniques with higher Recall@1 when $K$ is 50, 100, and 200, meaning it correctly ranks ground-truth of 3%, 1%, and 2% more queries as the most similar. Under the setting of $K=500$, Nova-1B ties with CodeArt with the same highest Recall@1. When looking into each individual benchmark, Nova-1B always wins on the most benchmarks under different settings of pool size $K$. For instance, Nova-1B wins on four benchmarks while DiEmph only wins on SQLite when $K=50$. We also conduct an ablation study on BCSD in Appendix A.5.
4.3 Analytic Experiments: How are Nova’s embeddings better?
@@
Compared with Nova-CL-HA (Figure 5 (a)), Nova-HA (Figure 5 (b)) including contrastive learning objectives in the pre-training, can separate the embeddings of assemblies with different functionalities better. Nova-HA clearly encode “Task 143” (orange points) away from the others. Yet, Nova’s (Figure 5 (c)) embeddings group the assemblies by functionalities more precisely than Nova-HA, suggesting that hierarchical attention enhances the training of contrastive learning objectives to learn more effective encoding. Visualization using a different approach, PCA, is shown in Appendix A.6. Analytic experiments on Nova’s hierarchical attention is shown in Appendix A.7.
Figure 5: t-SNE analysis of embeddings cauclated by Nova-CL-HA, Nova-HA, and Nova.
@@ 5.1 Binary Models
Machine learning models are widely used in binary program analysis tasks. However, these models are typically designed for specific tasks such as binary code similarity detection (Pei et al., 2020; Xu et al., 2023a; Wang et al., 2022), variable name prediction (Chen et al., 2022c; Xu et al., 2023b; Zhang et al., 2021), binary code type inference (Pei et al., 2021), and are built from scratch.
Recent techniques have started to pre-train foundation LLMs for binaries. CodeArt (Su et al., 2024) pre-trains encoder-style LLMs with a regularized attention design to better encode assembly code semantics. SLaDe (Armengol-Estapé et al., 2024) trains BART (Lewis et al., 2019) models on assembly. Meta LLMCompiler (Cummins et al., 2024) train CodeLlama models on LLVM IR to optimize binary code. LLM4Decompile (Tan et al., 2024) trains DeepSeekCoder with assembly for binary code decompilation. However, CodeArt does not generalize to generation tasks due to its encoder architecture. LLMCompiler trained on LLVM IR cannot be effectively transferred to assembly code. SLaDe and LLM4Decompile are limited in performance due to a lack of special designs for assembly. In contrast, Nova proposes hierarchical attention and contrastive learning objectives, outperforming existing techniques on both understanding (binary code similarity detection) and generation (binary code decompilation) tasks.
@@ 5.2 Large Source-Code Models
LLMs demonstrate promising results on many code-related tasks, such as code generation (Chen et al., 2022a; Liu et al., 2023; Chen et al., 2023; Le et al., 2022; Yue et al., 2021; Chen et al., 2021; Nijkamp et al., 2022; Fried et al., 2023; Rozière et al., 2023; Guo et al., 2024; Lozhkov et al., 2024; Hui et al., 2024), bug fixing (Jiang et al., 2023; Xia et al., 2023) and vulnerability fixing (Wu et al., 2023; Steenhoek et al., 2023; He & Vechev, 2023). The advances in using LLMs are attributed to the knowledge learned from massive source code and natural language text in their training datasets (Touvron et al., 2023; OpenAI, 2023). Nova is designed and trained for assembly, which has unique challenges such as low information density and diverse optimization.
5.3 Attention Mechanism
Standard self-attention is widely used in transformer architecture (Vaswani et al., 2023) to capture soft dependencies between tokens in the input. Many special attention mechanisms have been designed for better learning in various scenarios (Yang et al., 2016; Huang et al., 2024). LongCoder (Guo et al., 2023) combines window attention and global attention (attention sink (Xiao et al., 2024)) to handle long input of source code. We have shown that LongCoder’s window attention is less effective than Nova’s on assembly code.
CAST (Shi et al., 2021) is a new neural architecture that splits the abstract syntax tree (AST) of source code into subtrees, encodes the subtrees, and aggregates to the final encoding. PA-former (Chai & Li, 2024) is a new neural architecture that constructs source code as pyramid input based on their AST structure and contains a pyramid attention mechanism to calculate the features in a hierarchical way. HierarchyNet (Nguyen et al., 2023) is a neural architecture that considers source code AST, data flow, and control flow graphs. Similarly, it cannot be applied to assembly code. Different from CAST, PA-former, and HierarchiyNet, Nova’s attention design is for assembly code, is more lightweight and can be plugged into any pre-trained generative LLM.
6 Limitations
One limitation is that Nova is X86-specific, as we only collect X86 assembly corpus for pre-training. This design choice is mainly affected by two facts: (1) X86 assembly is used and explored in a wide range of binary tasks (Wang et al., 2022; Su et al., 2024; Xu et al., 2023a; Chen et al., 2022c) compared to other assembly languages, and (2) computation limitations. However, the proposed techniques are independent of X86 assembly. Low information density and compiler optimization are the common challenges of most assembly languages such as X86, ARM, and MIPS. The proposed techniques can be applied to the future development of multi-lingual assembly LLMs.